I keep getting 9 and 5/12
Stan drove 3 1/2 miles from home to work. He decided to go out for lunch and drove 1 3/4 miles each way to the local delicatessen. After work, he drove 1/2 mile to stop at the cleaners and then drove 3 2/3 miles home. How many miles did he drive in total?
the book is saying its 11 1/6
Thanks.

Found 2 solutions by stanbon, amoresroy:
Answer by stanbon(75887)

I keep getting 9 and 5/12
----------------------------------
Stan drove 3 1/2 miles from home to work. He decided to go out for lunch and drove 1 3/4 miles each way to the local delicatessen. After work, he drove 1/2 mile to stop at the cleaners and then drove 3 2/3 miles home. How many miles did he drive in total?
----------------------------------
(7/2) + (7/4) + (7/4) (1/2) + (11/3)
----
lcd = 12
----
[42 + 21 + 21 + 6 + 44]/12
----
= 134/12
---
= 67/6
---
= 11 1/6 miles
===================
Cheers,
Stan H.
===========================


the book is saying its 11 1/6

Answer by amoresroy(361)

I keep getting 9 and 5/12
Stan drove 3 1/2 miles from home to work. He decided to go out for lunch and drove 1 3/4 miles each way to the local delicatessen. After work, he drove 1/2 mile to stop at the cleaners and then drove 3 2/3 miles home. How many miles did he drive in total?
the book is saying its 11 1/6
3+6/12 + 1+9/12 + 1+9/12 + 6/12 + 3+8/12 = 11 1/6